OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. x2 − 16x + ____

jimthompson5910 (jim_thompson5910):

To find the missing term, take half of the x coefficient and then square it.

jimthompson5910 (jim_thompson5910):

Similar example: Find the missing term for x^2 + 8x + _____ Take half of the x coefficient 8 to get 8/2 = 4. Then square it to get 4^2 = 4*4 = 16 So the missing term is 16 This means x^2 + 8x + 16 is a perfect square trinomial.

OpenStudy (anonymous):

ok let me try brb

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

8^2

OpenStudy (anonymous):

then that would turn into 64?

jimthompson5910 (jim_thompson5910):

it's more like (-8)^2 = 64, but you have the correct final answer

OpenStudy (anonymous):

so x^2-16x+64?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ha thanks you made it a whole lot easier

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

ok but i have one more question just if you could explain it like the last one it is

OpenStudy (anonymous):

4x2 + ___x + 49

jimthompson5910 (jim_thompson5910):

This is slightly different, but follows the same idea

jimthompson5910 (jim_thompson5910):

If (a+b)^2 = 4x^2 + ___x + 49 Then we can FOIL out the left side to get (a+b)^2 = 4x^2 + ___x + 49 a^2 + ab + ab + b^2 = 4x^2 + ___x + 49 a^2 + 2ab + b^2 = 4x^2 + ___x + 49 So a^2 = 4x^2 2ab = ___x b^2 = 49 which means a = 2x 2ab = ___x b = 7 Since a = 2x and b = 7, this means 2ab = 2*2x*7 = 28x So 4x^2 + ___x + 49 becomes 4x^2 + 28x + 49

OpenStudy (anonymous):

wow ok let me see

jimthompson5910 (jim_thompson5910):

yeah its a lot

OpenStudy (anonymous):

what is the (a+b) the example or do i have to do it with everyone of these?

jimthompson5910 (jim_thompson5910):

(a+b)^2 is a perfect square and I used it to find the values of a and b...which helped me find 2ab (that middle term)

OpenStudy (anonymous):

ok